Circular No. 7624

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 HT_{50}

S. Pravdo, E. Helin and K. Lawrence, Jet Propulsion Laboratory, report the discovery of a possible comet by NEAT on CCD images obtained with the 1.2-m Schmidt at Palomar. T. B. Spahr, Minor Planet Center, has identified the object with the apparently asteroidal object 2001 HT_{50}, observed by LINEAR on Apr. 23 and by LONEOS on Apr. 26 (MPS 30375), and has now found observations back to Mar. 3. Further details and parabolic orbital elements (T = 2003 July 8, q = 2.80 AU, ω = 324^{o}, Ω = 43^{o}, i = 163^{o}, equinox 2000.0) are given on MPEC 2001-J31.

2001 UT α_{2000} δ_{2000} m_{2} Observer
Apr. 23.21809 11^{h}49^{m}06.^{s}20 −11^{o}25′18′′.3 18.4 LINEAR
26.13959 11 47 16.89 −11 12 36.1 18.5 LONEOS
May 14.22046 11 37 23.71 − 9 57 32.4 18.3 NEAT

SUPERNOVAE 2001bi, 2001bj, 2001bk, 2001bl, 2001bm, 2001bn, 2001bo

B. Vicente, K. Vieira, J. Cova, J. Hernández, C. Abad, and C. Briceño, Centro Investigaciones de Astronomia (CIDA), Merida, on behalf of the QUEST collaboration (cf. IAUC 7387), report on seven supernovae discovered with the QUEST 16-CCD array camera on the CIDA 1-m Schmidt telescope at Llano del Hato Observatory. The total area searched was 254 deg^{2} between Mar. 15 and 26, to R = 20.8, with each clear night receiving independent drift scans through B, V and R filters. Nightly scans continued until Apr. 4, so the seven objects all have well-sampled three-color lightcurves over roughly three weeks. Each supernova was found by subtraction of QUEST reference images taken between Mar. 3 and 8, and each new object has been confirmed via six or more independent images taken on three or more nights with subtractions from three independent reference images. The host galaxies for SNe 2001bj and 2001bl are not visible in QUEST data, to R ∼ 21.

SN 2001 UT α_{2000} δ_{2000} R Offset
2001bi Mar. 24  9^{h}28^{m}37.^{s}79 − 0^{o}36′58′′.1 20.0 1′′ W, 1′′ N
2001bj Mar. 24 10 36 25.83 − 0 58 17.8 20.2  —
2001bk Mar. 20 11 25 32.13 − 1 44 01.3 18.2 2′′ E, 4′′ S
2001bl Mar. 24 12 12 57.36 − 0 00 53.1 19.9  —
2001bm Mar. 25 13 35 58.65 − 0 45 03.3 19.7 2′′ E, 0′′ N
2001bn Mar. 25 14 30 39.53 − 2 06 00.6 19.4 1′′ W, 0′′ N
2001bo Mar. 25 15 25 40.80 − 1 18 00.5 20.8 1′′ W, 0′′ N
2001 May 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden